### v1

# Deep Neural Networks for Cross-Energy Particle Identification at RHIC and LHC

Abstract
This work demonstrates the application of a deep neural network for particle identification in the high transverse momentum regime. A model trained on simulated Large Hadron Collider (LHC) proton-proton collisions ($\sqrt{s}=13\,\mathrm{TeV}$) is used to classify nine distinct particles using seven kinematic-level features. The model is then tested on high transverse momentum RHIC data without any transfer learning, fine-tuning, or weight adjustment. It maintains accuracy above 91% for both LHC and RHIC sets, while achieving above 96% accuracy for all RHIC sets, including the $p_{T}>7\,\mathrm{GeV}/c$ set, despite not having been trained on any RHIC data. These results indicate that the model captures the underlying physics of high-energy collisions rather than just overfitting to the training data. This work highlights the potential of simulation-trained models to be deployed in different energy domains, especially in underrepresented or data-limited settings.
keywords: Deep learning, Particle identification, Generalization, Cross-domain, High $p_{T}$, RHIC, LHC
1 Introduction
Machine learning (ML) has played an increasingly central role in data collection and analysis in particle physics. Leading studies using deep neural networks have cemented machine learning as an essential tool at CERN’s LHC [1]. In particular, deep learning has shown significant promise in particle identification tasks ranging from jet classification [2], trigger systems, and calorimetry and event reconstruction [3]. Furthermore, it has become more critical in high momentum regimes where standard particle identification methods fail [4]. This large variety of machine learning applications is mainly applied to the Large Hadron Collider (LHC) data. In comparison, a lot less attention has been given to applying such methods in the context of the Relativistic Heavy Ion Collider (RHIC), despite its unique and rich physics environment.
Proton–proton collisions at RHIC, carried out at $\sqrt{s}=200\,\mathrm{GeV}$, represent a very different kinematic regime compared to that of the LHC’s $\sqrt{s}=13\,\mathrm{TeV}$. This difference in kinematics presents a valuable opportunity to test the ability of deep ML models trained on one center-of-mass energy to generalize to another. Such cross-regime generalization is rarely explored, and most ML efforts in high-energy physics are trained and tested on the same experimental or simulated data at the same energy scale.
In this study, we tested the ability of a simple deep neural network trained exclusively on LHC data to identify nine distinct but common particles in simulated proton-proton collisions. The model was trained on seven kinematic-level features and tested on various $p_{T}>3\,\mathrm{GeV}/c$ bins. Furthermore, we tested the model’s ability to generalize to RHIC’s lower center-of-mass energy under the same high $p_{T}$ conditions as the LHC set.
To our knowledge, this work represents one of the first attempts at LHC to RHIC generalization using deep learning, highlighting both the potential and the limitations of simulation-trained models in multi-regime applications. The findings suggest that with appropriate feature selection and network architecture, similar models can be used in real experimental settings, offering a practical solution when the training data is limited, noisy, or simply not available.
2 Dataset and feature design
PYTHIA 8[5] was used to simulate the p-p collision at both $\sqrt{s}=13\,\mathrm{TeV}$ and $\sqrt{s}=200\,\mathrm{GeV}$. There are two constraints on the generated data, a pseudorapidity cut $-2<\eta<2$ which corresponds to the central region of the detector, and a transverse momentum cut $p{{}_{{}_{T}}}>$ 3 GeV/c. The hadronic decay is turned on, which makes it necessary to choose particles that are long-lived and commonly produced in proton–proton collisions. This is because the model is not designed to deal with anomalies or particles that are quite rare and produced with low multiplicities. The particles selected are $\gamma$, $\pi^{+}$, $\pi^{-}$, $K^{+}$, $K^{-}$, $K_{L}^{0}$, $p$, $e^{-}$, $n$. The training set is a combination of two generated sets with slightly different parameters to facilitate data generation and event selection. It consists of a low $p_{T}$ set that has only the pseudorapidity condition and a high $p_{T}$ set with a $p_{T}>3\,\mathrm{GeV}/c$ cut. The training set contains 6.2 million events (5 million low $p_{T}$ and 1.2 million high $p_{T}$). Table 1 shows the generation parameters in detail for the training set and Table 2 shows the selected input features along with a brief description of them.
Table 1: Summary of Key PYTHIA8 Parameters Used in the Training Set
Parameter Name Low $p_{T}$ High $p_{T}$
Beams:idA 2212 (proton) 2212 (proton)
Beams:idB 2212 (proton) 2212 (proton)
Beams:eCM 13 TeV 13 TeV
Number of Events 5 million 1.2 million
PhaseSpace:pTHatMin 0 GeV 2.0 GeV
HardQCD:all on on
Charmonium:all off on
Bottomonium:all off on
PromptPhoton:all on on
Random:setSeed on on
PartonLevel:ISR on on
PartonLevel:FSR on on
HadronLevel:Decay on on
Table 2: Summary of the Input Features and Their Physical Significance
Input Feature Physical Significance Equation
Energy Energy from the relativistic energy-momentum relation $E^{2}=|\vec{p}|^{2}+m^{2}$
Momentum Magnitude of the momentum vector $|\vec{p}|=\sqrt{p_{x}^{2}+p_{y}^{2}+p_{z}^{2}}$
Transverse Momentum Magnitude of momentum in the transverse plane $p_{T}=\sqrt{p_{x}^{2}+p_{y}^{2}}$
Pseudorapidity Spatial angle relative to the beam axis $\eta=-\ln\left[\tan\left(\frac{\theta}{2}\right)\right]$
Rapidity Lorentz-invariant measure along the beam direction $y=\frac{1}{2}\ln\left(\frac{E+p_{z}}{E-p_{z}}\right)$
Azimuthal Angle Angle in the transverse plane around beam axis $\phi=\arctan\left(\frac{p_{y}}{p_{x}}\right)$
Charge Electric charge of the particle in units of $e$ $q$
3 Kinematic Feature Distributions and Class Composition
The input parameters are selected based on the criterion that they should be obtainable from laboratory measurements or derivable from other measured quantities. Moreover, the detectors that inspire the feature selection should be available at both the LHC and RHIC. Based on that information, the detectors responsible for the measurements would be hadronic and electromagnetic calorimeters[6, 7], time projection chambers[8, 9], and silicon vertex detectors. These detectors are foundational and essential for particle colliders. There is an additional constraint on the parameters, which is that they must be able to be generated using the event generator Pythia 8. The distributions of the key input features for both the low- and high-$p_{T}$ sets are shown in Figure 1, illustrating the kinematic differences between the two datasets.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 1: Summary of $p_{T}$, $\eta$, and $\phi$ distributions for low-$p_{T}$ and high-$p_{T}$ sets. The top row corresponds to the low-$p_{T}$ set and the bottom row corresponds to the high-$p_{T}$ set.
Table 3: Charge and particle distribution percentages in low and high transverse momentum training sets.
Particle Charge % in low $p_{T}$ in high $p_{T}$
$\gamma$ 0 52.193 11.198
$K_{L}^{0}$ 0 1.791 8.766
$n$ 0 1.016 6.578
$\pi^{+}$ +1 20.045 24.708
$p$ +1 1.024 6.580
$K^{+}$ +1 1.812 8.714
$e^{-}$ $-1$ 0.328 0.107
$K^{-}$ $-1$ 1.806 8.661
$\pi^{-}$ $-1$ 19.983 24.688
3.1 Momentum and Angular Distributions
[…]
The pseudorapidity distribution is symmetric around zero, as expected for particles produced in the central rapidity region in symmetric proton–proton collisions[11]. The $-2<\eta<2$ cut corresponds to a polar angle between $164.5^{\circ}$ and $15.5^{\circ}$, avoiding regions too close to the beamline. This distribution closely matches the rapidity distribution at high momentum under the condition that $p\gg m$. The azimuthal distribution is completely symmetric and uniform, indicating isotropic particle production. Moreover, the values of the azimuthal angle span the full range $[-\pi,\pi]$, providing full coverage around the beam axis.
3.2 Low-$p_{T}$ Particle Composition
In the low $p_{T}$ regime shown in Table 3, the most dominant particles are neutral particles, which is natural since approximately $52\%$ of the particles are photons. These photons are primarily produced from the decay of neutral pions via $\pi^{0}\rightarrow 2\gamma$. This decay has a branching ratio of approximately $98.8\%$ [12], and is a major contributor to the photon abundance. Additionally, photons can be produced directly during collisions or as byproducts of other decays.
The second most abundant particles are the positive and negative pions, each constituting about $20\%$ of the particles. Their abundance is due to their low mass, making them kinematically favored in QCD interactions. Their numbers are equal due to isospin symmetry, as $\pi^{+}$, $\pi^{0}$, and $\pi^{-}$ form an isospin triplet [13]. Had the neutral pion not decayed promptly, the number of $\pi^{0}$ would have matched the $\pi^{\pm}$ counts, but its rapid decay into two photons effectively boosts the photon yield relative to charged pions.
Protons and neutrons, which form an isospin doublet [13], are produced in approximately equal numbers due to having similar mass and quark structure. Their overall lower production rate compared to pions is attributed to their higher mass, making their production less kinematically favorable.
Kaons, which include a strange quark or antiquark, are produced in approximately equal numbers between $K^{+}$ and $K^{-}$ as a reflection of strangeness conservation. However, strange quarks are subject to strangeness suppression [14], leading to fewer kaons compared to pions. The kaon-to-pion ratio ($K/\pi$) calculated as $0.135$ is consistent with typical values observed in high-energy proton-proton collisions [15].
Finally, electrons represent the least abundant particle type because they are produced via electromagnetic processes, which are less dominant compared to strong interactions in these collisions.
3.3 High-$p_{T}$ Behavior and Transitions
[…]
The particle type and charge distributions exhibit notable shifts at high transverse momentum. Photons, previously dominant at low $p_{T}$, now account for only about $11\%$ of the total. While the $\pi^{0}\rightarrow 2\gamma$ decay process still occurs, the resulting photons often fall below the $p_{T}>3\,\mathrm{GeV}/c$ threshold. Consequently, the charged pion contribution rises, with $\pi^{+}$ and $\pi^{-}$ together making up about $50\%$ of the particles.
The kaons experience a significant boost in their production, rising from approximately $1.8\%$ to $8.7\%$ for both $K^{+}$ and $K^{-}$. This increase is attributed to the greater availability of energy in collisions, partially overcoming strangeness suppression. Similarly, protons and neutrons also exhibit increased production due to the dominance of hard scattering and fragmentation processes at higher energies. Even at high $p_{T}$, electrons still constitute only about $0.1\%$ of all produced particles, again highlighting the influence of strong interaction processes.
3.4 Suitability for Model Training
All previously examined kinematic features and distributions are consistent with theoretical expectations and align well with experimental observations from proton-proton collisions. The transition from a photon-dominated distribution to a more dynamic and energetic particle composition reflects the increasing influence of hard scattering and fragmentation processes. The $|\eta|<2$ cut focuses the analysis on the central detector region, where particle identification is most effective. Combining soft particles from the low momentum set with rarer, high-momentum particles ensures coverage of a broad range of the kinematic spectrum of proton-proton collisions. The physical validity and diversity of the combined dataset make it highly suitable for training the model to perform particle identification across a wide momentum range.
4 Neural Network Training
The neural network architecture chosen is a standard feed-forward network, specifically a multi-layer perceptron (MLP)[16]. The hidden layers use the Rectified Linear Unit (ReLU) activation function, a computationally efficient nonlinear activation function defined as[17]
$f(x)=\max(0,x).$ (1)
Compared to other alternatives such as tanh or sigmoid, ReLU avoids exponentiation. This makes it suitable for handling large datasets such as those from LHC simulations.Non-linearity in hidden layers in essential in capturing complex relationships between inputs and outputs.
The output layer activation function and the loss function are closely related and therefore are discussed together. The SoftMax activation function is a well suited for multi-class classification which is the goal of the network. The function converts the values it receives from the hidden layer into a probability distribution of the classes, with the sum of probabilities being equal to one. The SoftMax function takes the form: [18]
$f(x_{i})=\frac{e^{x_{i}}}{\sum_{j}e^{x_{j}}}$ (2)
where $f(x_{i})$ is the predicted probability, the denominator normalizes the output so that the sum of probabilities is equal to one. This predicted class corresponds to the highest output probability.
The network is trained to minimize the categorical cross-entropy loss, defined as [19]
$\text{CCE}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{j=1}^{c}y_{ij}\log\left(\hat{y}_{ij}\right)$ (3)
where $y_{ij}$ represents the true label of the class while $\hat{y}_{ij}$ represents the predicted probability of the class (output of the SoftMax function). $N$ is the total number of samples and $c$ is the number of classes. This function measures how close the predicted distribution is to the true distribution of class labels.
The optimizer is responsible for updating the weights of the model during training by adjusting key parameters like the learning rate. For this current setup, the Adaptive Moment Estimation (ADAM) optimizer is used [20]. It is adaptive as it adjusts the individual learning rates of each parameter. It combines the benefits of Root Mean Square Propagation(RMSProp)[21], including adaptive learning rates, with additional advantages such as bias correction and momentum, making it well-suited for the network at hand.
For the network’s structure, each input parameter is assigned to one neuron in the input layer; therefore, the input layer has seven neurons. The number of hidden layers depends on the complexity of the task, but having two or more makes the network a deep neural network. Deep neural networks are suitable for tasks where standard methods struggle, due to their ability to capture complex patterns. The number of neurons in the hidden layer has a wide range of integer values depending on the application. However, a good starting point is having an equal number of neurons as the input layer. This network underwent many different iterations before settling on the final structure. The initial structure was 7 neurons in both hidden layers; however, to reduce overfitting, the number of neurons was reduced to 5 in the second hidden layer. Finally, the output layer uses a SoftMax activation function, where each neuron represents the probability of a class. Since there are nine particle types, the output layer contains nine neurons. Figure 2 shows a schematic representation of the network.
Figure 2: Structure of the deep neural network
The network was trained for 34 epochs with a batch size of 128, taking an average of 41.5 seconds per epoch and about 23.5 minutes to train. The chosen regularization method is early stopping[22] with a patience value of 8. The model continues iterating until there are no improvements in validation loss for 8 consecutive epochs. Afterwards, the model restores the weights from the epoch with the lowest validation loss. This training technique improves in generalization and reduces overfitting on future test sets.
The dataset is split into three parts: training, validation, and test sets. The training set is used by the model to learn the relationships and patterns between inputs and outputs. The model iterates over the training data for 34 epochs. During training, the model receives the inputs and the true label or identification of the particle to learn the mapping between the input parameters and the output labels.
The validation consists of examples which the model hasn’t seen. It is used after each epoch to check overall how it generalizes to unseen data. The validation set is crucial in guiding and improving the performance of the model and avoiding issues such as overfitting, where the model performs well in training but fails generalizing to new examples.
The test set is used to assess the overall performance of the model after applying all the improvements and tuning during training. It is only used once, after the model has finished training, to examine how the model performs on real unseen data.
The dataset is split into an 80% training, 10% validation, and 10% test sets. Although the 90-5-5 split is common, the 80-10-10 split was chosen to maintain a large number of examples, in this case about 5 million training examples while still having a considerable validation and test set of 620 thousand examples. Figure 3 shows the change in validation loss during training. The point of minimum loss is highlighted by a star.
Figure 3: Validation loss vs. epochs
For the normalization of the training data, by observing the values of all the input parameters, they all fall within manageable ranges where the minimum is $-3.14$ (from the azimuthal angle). In terms of transverse momentum, most particles fall within a range of 00 to $50\,\mathrm{GeV}/c$, with a small number of outliers having $100$ to $200\,\mathrm{GeV}/c$. Naturally, for the low momentum set without the $p_{T}>3\,\mathrm{GeV}/c$ cut, most of the generated particles are soft particles having a transverse momentum value between 00 and $2\,\mathrm{GeV}/c$. Since the values for all input parameters are within close range of each other, normalization was deemed to be unnecessary.
Table 4: Summary of the training, validation, and test set performances of the neural network.
Point of Comparison Training Set Validation Set Test Set
Accuracy (%) 99.47 99.88 99.88
Loss 0.0172 0.0061 0.0058
Number of examples 4,960,000 620,000 620,000
From Table 4, the model performs well on all three sets, having comparable accuracy above 99%. The lowest accuracy among the three is the training set, but that can be attributed to its large size while the validation and test sets represent only a small subset of the total data. This means the training set contains a broader range of high momentum events, which explains its slightly reduced accuracy.
Overall, the model has no issues generalizing to unseen data based on the high performance and low loss values for both validation and test sets. The model is subsequently tested on specific cuts of high transverse momentum subsets at both RHIC and LHC energies.
5 Test Sets Results
The model was tested on both RHIC and LHC sets at a transverse momentum greater than $3\,\mathrm{GeV}/c$. The high-$p_{T}$ range was probed further by dividing them into $p_{T}$ bins with $3$-$5$, $5$-$7$, and $>7\,\mathrm{GeV}/c$. The data were generated with the same conditions as the high-$p_{T}$ of the training data in Table 1. The only change was setting pTHatMin=4.0 GeV for convenience and to facilitate event generation. The model was tested specifically on RHIC to verify its ability to generalize to lower center-of-mass energy, as it was trained on LHC’s $\sqrt{s}=13\,\mathrm{TeV}$. The performance results are summarized in Table 5.
Table 5: Summary of the model’s performance on the RHIC and LHC test sets across different $p_{T}$ ranges.
Point of Comparison High $p_{T}$ LHC High $p_{T}$ RHIC 3-5 LHC 3-5 RHIC 5-7 LHC 5-7 RHIC 7+ LHC 7+ RHIC
Accuracy (%) 99.44 99.70 99.80 99.87 98.24 98.66 91.33 96.23
Loss 0.0291 0.0180 0.0181 0.0129 0.0716 0.0467 0.2380 0.1108
Number of examples 500,000 500,000 50,000 50,000 50,000 50,000 50,000 50,000
The model achieves high performance in the high transverse momentum regions, retaining an accuracy above 91% for all test sets at both LHC and RHIC. Moreover, it performs as expected, having slightly lower accuracy as it moves into higher transverse momentum regions.
6 Discussion
The key difference between the two sets is strikingly visible at the $p_{T}>7\,\mathrm{GeV}/c$ cut, where the RHIC set has much better accuracy at 96.23% compared to 91.33% of the LHC set. This difference can be explained by the training of the model and the nature of the LHC set. The model has been trained on the LHC set at $\sqrt{s}=13\,\mathrm{TeV}$, which means that the particles are much more energetic overall, with momentum ranges that are much higher than RHIC, especially at the transverse momentum range of $p_{T}>7\,\mathrm{GeV}/c$. The performance of the model on different center of mass energies can be better understood by comparing the transverse momentum distributions in Figure 4.
A dataset with the condition of $p_{T}>3\,\mathrm{GeV}/c$ includes a lot of high momentum particles, but relatively few in the $p_{T}>7\,\mathrm{GeV}/c$ region, which contributes to the drop in accuracy for LHC. For any generated set, most of the particles produced have the lowest possible energy and momentum, which corresponds to $3\,\mathrm{GeV}/c$ for the training set and $7\,\mathrm{GeV}/c$ for test set.
Refer to caption Refer to caption
(a) LHC $p_{T}$ distribution (b) RHIC $p_{T}$ distribution
Figure 4: Comparison of transverse momentum ($p_{T}$) distributions between (a) LHC $\sqrt{s}=13\,\mathrm{TeV}$ proton-proton collisions and (b) RHIC $\sqrt{s}=200\,\mathrm{GeV}$ proton-proton collisions with $p_{T}>3\,\mathrm{GeV}/c$ cut. Both distributions are shown on a logarithmic scale to highlight the wide dynamic range of particle momenta. The LHC distribution shows a broader and harder $p_{T}$ spectrum compared to RHIC. On the other hand, the $p_{T}$ distribution at $\sqrt{s}=200\,\mathrm{GeV}$ is noticeably softer with a much steeper fall off as measured by the PHENIX Collaboration [23].
7 Conclusion
This paper demonstrates the ability of a fundamentally simple deep neural network trained exclusively on simulated LHC proton-proton collisions ($\sqrt{s}=13\,\mathrm{TeV}$), to generalize to RHIC data at a much lower center-of-mass energy. This generalization is achieved without any fine-tuning, weight adjustment or transfer learning. The model maintains an accuracy above 91% at all momentum cuts, even in the complex $p_{T}>7\,\mathrm{GeV}/c$ region. For the RHIC set under the same conditions, it achieves an accuracy of 96.23%, which is quite notable as the model was not trained on any $\sqrt{s}=200\,\mathrm{GeV}$ examples. The observed difference in the model’s performance is consistent with expectations, given the higher kinematic reach of the LHC. The ability of the model to generalize seamlessly to a lower center-of-mass energy highlights its robustness and suggests that it captures the underlying physics rather than overfitting to the kinematics of each regime.
These findings open the door to applying experimental simulation models in different energy domains beyond their original ones, especially in scenarios where training data is limited or simply unavailable. Future work will focus on introducing more detector-level features and employing more sophisticated network architectures, such as combined networks to handle the noisy nature of real experimental data.